# Extracting Signal Electron Trajectories in the COMET Phase-I Cylindrical Drift Chamber Using Deep Learning

### v1

@@
We present a pioneering approach to tracking analysis within the COMET Phase-I experiment, which aims to search for the charged lepton flavor violating $\mu\to e$ conversion process in a muonic atom, at J-PARC, Japan. This paper specifically introduces the extraction of signal electron trajectories in the COMET Phase-I cylindrical drift chamber (CDC) amidst a high background hit rate, with more than 40 % occupancy of the total CDC cells, utilizing deep learning techniques of semantic segmentation. Our model achieved remarkable results, with a purity rate of 98 % and a retention rate of 90 % for CDC cells with signal hits, surpassing the design-goal performance of 90 % for both metrics. This study marks the initial application of deep learning to COMET tracking, paving the way for more advanced techniques in future research.
@@ 1 Introduction
The COMET experiment is to search for charged lepton flavor violation (CLFV), with a specific emphasis on neutrinoless coherent muon to electron conversion ($\mu\to e$ conversion) process in a muonic atom, namely,
@@
where $N(A,Z)$ denotes a nucleus with mass number $A$ and atomic number $Z$. This CLFV process is considered to be a promising avenue for discovering new physics beyond the Standard Model (SM)[1]. In the minimal SM, neutrinos were thought to be massless, and lepton flavor conservation was believed to hold within each generation. However, neutrino oscillation experiments revealed that neutrinos have mass and can mix flavors, disproving lepton flavor conservation. However, the SM predicts an extremely small contribution to CLFV rates, whose contribution is astonishingly small, approximately at the rate of the order of $10^{-54}$, offering a unique experimental avenue to explore new physics beyond the SM (BSM) without significant interference from SM backgrounds. Expected sensitivities of the upcoming CLFV experiments, including the COMET experiment, enable the exploration of new physics up to energy scales of O($10^{4}$) TeV, surpassing the capabilities of collider experiments. Thus, CLFV research becomes a good complement to Beyond the Standard Model (BSM) searches at the Large Hadron Collider.
The COMET experiment will be conducted at the Japan Proton Accelerator Research Complex (J-PARC), Tokai, Japan. The COMET stands for Coherent Muon to Electron Transition. It will be proceeded in a staged approach, where Phase-I of the COMET experiment will aim at an experimental sensitivity of $7\times 10^{-15}$ with 90 % confidence level. It is about 100 times improvement over that latest result of [2]. The Phase-II, the second stage of the COMET experiment, would aim at further improvement of experimental sensitivity of the order of $\mathcal{O}(10^{-17})$. Further refinement of additional factor of ten, down to $\mathcal{O}(10^{-18}$) has been attempted.
@@
Figure 2 illustrates the COMET Phase-I detector layout, featuring the Cylindrical Drift Chamber (CDC). Being placed after the muon transport section, the CDC resides within the warm bore of a large 1 T superconducting detector solenoid. Additionally, Cylindrical Trigger Hodoscopes (CTH), consisting of two layers of plastic scintillation counters, are placed upstream and downstream on the inner side of the CDC to trigger events and provide reference timing of the events. The CDC plays an important role in the primary momentum measurement for electrons resulting from $\mu\to e$ conversion, crucial for identifying and measuring signal electrons while efficiently rejecting backgrounds.
@@ 2 COMET Tracking Analysis
The physics analysis will focus on identifying signal events and assessing background contributions. The identification of signal events from $\mu\to e$ conversion is primarily based on their momentum determination. For an aluminium target, which will be used in the COMET Phase-I experiment, the momentum of the signal events is 104.9 MeV/c. The momentum determination by tracking at the CDC is of critical importance in the COMET Phase-I experiment.
The challenge facing CDC tracking lies in effectively distinguishing signal hits from numerous background hits, originating mainly from beam flashes and products resulting from muon decays and muon nuclear capture, in the CDC. With the world’s highest muon beam intensity employed in COMET Phase-I of $\mathcal{O}(10^{9})$ stopped muons/s, being over 10 times greater than that at Paul Scherrer Institute (PSI), Switzerland, the scale of background hits is potentially substantial. From the studies of this issue with simulation data, an anticipated background hit occupancy of approximately 40 % or more for all active cells in the CDC emerges. In addition, potential fluctuations in intensity of each pulsed proton beam bunch by a factor of two or more has been identified in the past proton beam studies at J-PARC. This emphasises the critical need to assess how efficiently background hits can be rejected while retaining as many signal hits as possible. The success of CDC tracking performance depends on optimizing this balance, addressing the complexities posed by the unprecedented intensity of muon beams in the COMET Phase-I experiment.
Previously, the exploration of signal hit extraction methods has included the use of gradient-boosted decision trees (GBDT) and the application of the Hough transformation, both of which are valuable for detecting circles amidst significant background hits. These approaches have demonstrated good efficiency up to a background hit occupancy of approximately 15 %. [3]. However, the increasing uncertainty regarding the actual background hit occupancy, coupled with potential fluctuations in proton intensity by a factor of two or more, requires a quest for a more robust and resilient method.
@@
For effective hit filtering, various features of the hits are utilized. These include the spatial location of the cells with hits and specific characteristics of each hit recorded by the CDC. The CDC records timestamps of incoming pulses (hit times) relative to the CTH reference timing and the sizes of these pulses (hit ADCs), which reflect the charge deposited by the hits. Additionally, the relative location of the cell to the CTH trigger counter is valuable information, as the track spans a limited spatial region close to the CTH trigger counter. This azimuthal angle information from the CTH trigger counter, referred to as $\Delta\phi$, is utilized.
@@ 3.1 Beam and Detector Simulations
The full simulation of the muon beam begins with protons hitting the proton target to produce pions. This was carried out using the standard software framework (ICEDUST [3]), which utilizes Geant4. The simulation included the complete geometry of the experimental hall and a comprehensive magnetic field map, taking fringe fields into account. In Figure 1 a display of a small fraction of the muon beam in this simulation is shown.
After simulating the muon beam, a simplified detector response simulation was performed to model the response of the working gas in the CDC, based on Garfield simulation results. In this step, waveforms from the CDC readout electronics were generated based on the incident particles recorded in the CDC from the aforementioned muon beam simulation. The waveforms were then converted into pairs of ADC and time, which we refer to as hits hereafter.
@@
Eventually, the hits sequence ring was divided into 300 event frames to examine the background hits distribution. The hit occupancy distribution is around 40 %, as shown in Figure 4. The 2-dimensional distribution of charge and timing for all the background hits are shown in Figure 5. It should be noted that there are sometimes multiple hits in one cell from a single event. The multiplicity distribution according to this simulation is shown in Figure 6
@@ 3.2 Background Hit Samples from Resampling
Due to the limited computing resources, it is not practical to directly generate millions of events for our study using the full simulation method. Instead, we employed a resampling method based on the distributions mentioned above to create a sufficient number of background hit samples.
It is important to note that this method cannot account for the correlation among individual hits. As a result, some cluster patterns of the background hits in the CDC, which are observable from the full beam simulation, are not preserved in the samples generated using the resampling method. The performance, considering cluster patterns, will be studied in future work once we find an appropriate way to generate background hit samples with this feature.
@@ 3.3 Signal Electron Samples
After the selection, a pure signal sample was generated. The typical number of signal hits per signal electron is about 66. The 2-D distribution of all the signal hits is shown in Figure 7.
@@ 3.5.1 Cells with Multiple Hits
This approach still requires using one hit to determine the features of a cell. We employed two methods to select which hit to use. The first method uses the hit time and hit ADC distribution of signal hits, as shown in Figure 7. In this method, we create a 2D histogram where each axis represents a different feature: hit time and ADC value. This histogram is populated with signal hits from the simulation, creating a density map. The density at each point in this 2D distribution indicates the likelihood that a hit with those particular time and ADC values is a signal hit. For a cell with multiple hits, we examine the time and ADC values of each hit and select the hit that corresponds to the highest density (or likelihood) in the 2D histogram. By selecting the hit with the highest likelihood, we can achieve reasonable prediction accuracy.
For comparison, we also tested a simpler method of selecting the first hit in each cell. Fortunately, both methods showed similar performance on the small sample from the full simulation and the large sample from the resampling method.
@@ 3.5.2 Converting Hits to Pixels
After reducing the dimensionality by selecting only one hit per cell, each pixel corresponds to a single hit. Thus, only three color channels are needed for pixels. The first channel represents the hit time, the second channel represents the ADC sum, and the third channel represents $\Delta\phi$, the azimuthal angle between the CTH trigger and each cell.
The three types of input values can be converted to color values using simple linear scaling. However, one feature in the distribution of signal hits suggests that scaling with a limited input value range might be beneficial for maintaining precision with a limited number of bits for color values. In the simulation, as shown in Figure 8, most signal electron hits fall within the range (--50.0 ns, 400 ns) for time information and (0, 700) for the ADC information. In the limited scaling scheme, hits within these ranges are linearly scaled, while those outside are assigned cap values.
When mapping hit cells to pixels, some pixels correspond to non-existing cells at all, while some correspond to cells with no hits. To distinguish these special pixels from others, black is reserved for pixels with no corresponding cells, and white is used for pixels corresponding to empty cells. Narrow gaps in the color value range are maintained around both black and white to avoid confusion between these special cases and regular hit information.
Figure 8: Signal and background hit distributions for hit time (left) and hit ADC (right) with log-scaled y-axis. Signal hits cluster in specific ranges, while background hits are uniformly distributed.
@@ 3.5.3 Arranging Pixels on Canvas
In the ”normal” format, the detector’s 2D plane—more specifically, a projection of cells at the CDC readout plane—is treated as an image, using (x, y) coordinates to define the plane on the CDC section. The cell size is approximately 16.8 mm wide and 16.0 mm high, with the radius of the CDC readout at 818 mm [3]. To represent each detector cell as a single pixel, the plane is discretized into 128 bins along both the x and y axes, resulting in a 128$\times$128 pixel image. This results in a grid size of 12.8 mm per bin (1636 mm / 128), which is smaller than the actual cell size. Although the cells are arranged along the section of a cylinder, requiring a finer grid when using Cartesian coordinates, this discretization is essential for ensuring each cell is accurately represented by a single pixel. However, it is important to note that approximately 70% of this image area remains unoccupied(black area of Figure 9(a)), rendering this format inefficient for neural network training.
@@
Figure 9: (a),(b) The input image and label of the normal format. (c),(d) The input image and label of the reduced format. Each input image has 3 channels: red for hit time, green for ADC of hits, and blue for $\Delta\phi$ of hits. White pixels of input image indicate cells with no hits, while black pixels represent areas not corresponding to any cells in the CDC detector. The label images are grayscale, with white pixels denoting cells with true signal hits.
@@ 4.1.1 U-Net
EfficientNet, developed through neural architecture search [20, 21], optimizes both accuracy and efficiency by scaling up depth, width, and resolution in a balanced manner. This model builds on the success of earlier architectures like ResNet [4], incorporating advancements that allow it to achieve superior performance with fewer parameters and reduced computational cost. EfficientNet comes in a family of models ranging from EfficientNet-B0 to EfficientNet-B7, each with an increasing number of parameters and layers. This scalability allows us to evaluate different model capacities by simply changing the model index, facilitating a more systematic investigation of the model’s performance on our task.
@@ 5 Implementation Details
Our preprocessing methods include options such as linear scaling and limited scaling for the scaling method, and first hit and 2D histogram for multi-hit handling. By default, we use linear scaling for the scaling method and the 2D histogram method for multi-hit handling.
@@ 6 Results and Analysis
This metric provide a comprehensive evaluation of the model’s ability to accurately identify positive and negative cases, a critical aspect for addressing the requirements of COMET tracking task. And it is said that to obtain the necessary energy resolution in COMET, this Purity-Retention curve must pass outside (0.9,0.9), so we will discuss whether such a learning result can be obtained.
@@ 6.1 Model Performance on Signal Detection
Our model achieved the target score, surpassing both purity and retention metrics of 0.9 simultaneously, as illustrated in Figure 11. We can visually confirm our model successfully picks up target signal information from noisy input as shown in Figure 12. From Figure 11, we confirmed two important points to achieve the target score:
@@ 6.2 Effect of Scaling Model Parameters
Another reason for the less sensitivity of parameter scaling at larger image sizes is the nature of the current background hits, which are generated with the resampling method (Sec.3.2) and can be considered relatively easy to classify compared to real cases. Consequently, the models can achieve high accuracy even with fewer parameters when the image size is sufficient to capture the necessary details.
@@ 6.3 Handling Multiple Hits
The Purity-Retention curves in Figure 15 compare the performance of two pre-processing methods for handling multiple hits within the same cell: the naive 1st-hit method and the 2D histogram method. The left curve shows the overall Purity-Retention relationship, while the right zooms in on the high-purity region (close to 1.0). The 2D histogram method (green dashed line) consistently outperforms the 1st-hit method (blue solid line) across almost the entire range of retention values, particularly in the high-retention region, where it maintains higher purity levels compared to the 1st-hit method.
The superior performance of the 2D histogram method can be attributed to its enhanced signal hit coverage, as highlighted in Table 3. The table shows that the 1st-hit method has lower overall signal coverage (87%) and significantly lower signal coverage for multi-hit cells (74%). In contrast, the 2D histogram method achieves a higher overall signal coverage (97%) and better coverage for multi-hit cells (94%). This improved coverage translates directly into higher purity at comparable retention levels, as demonstrated in the Purity-Retention curves. Although the techniques we explored are relatively basic, their promising results warrant further exploration in future studies.
Figure 15: Purity-Retention curves for two methods. The left plot shows the full-scale view of the Purity-Retention curves for the 1st_hit (solid blue line) and the highest_hit (dashed green line). The 1st_hit refers to using the first hit method for multiple hits handling, while the 2D_histogram refers to selecting the highest value with the 2D-histogram method to improve model performance. The right plot is a zoomed-in view focusing on the high-purity and high-retention region, indicated by the yellow area. This highlighted area represents where both purity and retention are above 0.9, which is our target.
Table 3: Comparison of Signal Coverage between 1st Hit and 2D Histogram methods. This statistics is calculated with 8000 events from the training data. The 2D histogram method shows better coverage than the 1st hit method, which explains the superior performance of the 2D histogram Purity-Retention curve.

### v2

@@
We present a pioneering approach to tracking analysis within the COMET Phase-I experiment, which aims to search for the charged lepton flavor violating $\mu\to e$ conversion process in a muonic atom, at J-PARC, Japan. This paper specifically introduces the extraction of signal electron trajectories in the COMET Phase-I cylindrical drift chamber (CDC) amid a high background hit rate, with more than 40 % occupancy of the total CDC cells, utilizing deep learning techniques of semantic segmentation. Our model achieved remarkable results, with a purity rate of 98 % and a retention rate of 90 % for CDC cells with signal hits, surpassing the design-goal performance of 90 % for both metrics. This study marks the initial application of deep learning to COMET tracking, paving the way for more advanced techniques in future research.
@@ 1 Introduction
The COMET experiment aims to search for charged lepton flavor violation (CLFV), with a specific emphasis on neutrinoless coherent muon to electron conversion ($\mu\to e$ conversion) process in a muonic atom, namely,
@@
where $N(A,Z)$ denotes a nucleus with mass number $A$ and atomic number $Z$. This CLFV process is considered to be a promising avenue for discovering new physics beyond the Standard Model (SM)[1]. In the minimal SM, neutrinos were thought to be massless, and lepton flavor conservation was believed to hold within each generation. However, neutrino oscillation experiments revealed that neutrinos have mass and can change flavors, disproving the assumption of strict lepton flavor conservation. However, the SM predicts an extremely small contribution to CLFV, with a rate on the order of $10^{-54}$, offering a unique experimental avenue to explore new physics beyond the SM (BSM) without significant interference from SM backgrounds. Expected sensitivities of the upcoming CLFV experiments, including the COMET experiment, will allow the exploration of new physics up to energy scales of O($10^{4}$) TeV, surpassing the capabilities of collider experiments. Thus, CLFV research serves as an important complement to Beyond the Standard Model (BSM) searches at the Large Hadron Collider.
The COMET experiment will be conducted at the Japan Proton Accelerator Research Complex (J-PARC), Tokai, Japan. COMET stands for Coherent Muon to Electron Transition. The experiment will proceed in a staged approach, where Phase-I of the COMET experiment will aim at an experimental sensitivity of $7\times 10^{-15}$ at a 90 % confidence level. It is about 100 times improvement over that latest result of [2]. Phase-II, the second stage of the COMET experiment, will aim at further improvements in sensitivity, down to the order of $\mathcal{O}(10^{-17})$. A refinement of additional factor of ten, down to $\mathcal{O}(10^{-18}$) has been attempted.
@@
Figure 2 illustrates the COMET Phase-I detector layout, featuring the Cylindrical Drift Chamber (CDC). Being placed after the muon transport section, the CDC resides within the warm bore of a large 1 T superconducting detector solenoid. Additionally, Cylindrical Trigger Hodoscopes (CTH), consisting of two layers of plastic scintillation counters, are placed upstream and downstream on the inner side of the CDC to trigger events and provide reference timing of the events. The CDC is essential for measuring the momentum of electrons from $\mu\to e$ conversion, enabling accurate identification of signal electrons and effective rejection of background events.
@@ 2 COMET Tracking Analysis
The physics analysis focuses on identifying signal events and evaluating background contributions. For $\mu\to e$ conversion, signal events are primarily identified by their momentum. In the COMET Phase-I experiment with an aluminum target, the signal momentum is 104.9 MeV/c. Accurate momentum tracking in the CDC is therefore crucial to the success of COMET Phase-I.
A major challenge in CDC tracking lies in effectively distinguishing signal hits from numerous background hits, originating mainly from beam flashes and products resulting from muon decays and muon nuclear capture at the muon target. With the world’s highest muon beam intensity employed in COMET Phase-I of $\mathcal{O}(10^{9})$ stopped muons/s, being over 10 times greater than that at Paul Scherrer Institute (PSI), Switzerland, the scale of background hits is potentially substantial. From the studies of this issue with simulation data, an anticipated background hit occupancy of approximately 40 % or more for all active cells in the CDC emerges. In addition, potential fluctuations in intensity of each pulsed proton beam bunch by a factor of two or more has been identified in the past proton beam studies at J-PARC. This emphasises the critical need to assess how efficiently background hits can be rejected while retaining as many signal hits as possible. The success of CDC tracking performance depends on optimizing this balance, addressing the complexities posed by the unprecedented intensity of muon beams in the COMET Phase-I experiment.
Previously, the exploration of signal hit extraction methods has included the use of gradient-boosted decision trees (GBDT) and the application of the Hough transformation, both of which are valuable for detecting circles amid significant background hits. These approaches have demonstrated good efficiency up to a background hit occupancy of approximately 15 %. [3]. However, the increasing uncertainty regarding the actual background hit occupancy, coupled with potential fluctuations in proton intensity by a factor of two or more, requires a quest for a more robust and resilient method.
@@
For effective hit filtering, various features of the hits are utilized. These include the spatial location of the cells with hits and specific characteristics of each hit recorded by the CDC. The CDC records timestamps of incoming pulses (hit timing) relative to the CTH reference timing and the sizes of these pulses (hit ADCs), which reflect the charge deposited by the hits. Additionally, the relative location of the cell to the CTH trigger counter is valuable information, as the track spans a limited spatial region close to the CTH trigger counter. This azimuthal angle information from the CTH trigger counter, referred to as $\Delta\phi$, is utilized.
@@ 3.1 Beam and Detector Simulations
The full simulation of the muon beam begins with protons hitting the proton target to produce pions. This was carried out using the COMET standard software framework (ICEDUST [3]), which utilizes Geant4. The simulation included the complete geometry of the experimental hall and a comprehensive magnetic field map, taking fringe fields into account. In Figure 1 a display of a small fraction of the muon beam in this simulation is shown.
After simulating the muon beam, a simplified detector response simulation was performed to model the response of the working gas in the CDC, based on Garfield simulation results. In this step, waveforms from the CDC readout electronics were generated based on the incident particles recorded in the CDC from the aforementioned muon beam simulation. The waveforms were then converted into pairs of ADC and timing, which we refer to as hits hereafter.
@@
Eventually, the hits sequence ring was divided into 300 event frames to examine the background hits distribution. The average hit occupancy across the full time range is approximately 40 %, as shown in Figure 4. The 2-dimensional (2D) distribution of charge and timing for all the background hits are shown in Figure 5. It should be noted that there are sometimes multiple hits in one cell from a single event. The multiplicity distribution according to this simulation is shown in Figure 6
@@ 3.2 Background Hit Samples from Resampling
Due to the limited computing resources, it is not practical to directly generate millions of events for our study using the full simulation method. Instead, we employed a resampling method based on the distributions mentioned above to create a sufficient number of background hit samples. It is important to note that this method cannot account for the correlation among individual hits. As a result, some cluster patterns of the background hits in the CDC, which are observable from the full beam simulation, are not preserved in the samples generated using the resampling method. The performance, considering cluster patterns, will be studied in future work.
@@ 3.3 Signal Electron Samples
After the selection, a pure signal sample was generated. The typical number of signal hits per signal electron is about 66. The 2D distribution of timing versus charge for all signal hits is shown in Figure 7.
@@ 3.5.1 Cells with Multiple Hits
This approach still requires using one hit to determine the features of a cell. We employed two methods to select which hit to use. The first method uses the hit timing and hit ADC distribution of signal hits, as shown in Figure 7. In this method, a 2D distribution is created with each axis representing a distinct feature: hit timing and ADC value. This distribution is populated with signal hits from the simulation, creating a density map. The density at each point in this 2D distribution indicates the likelihood that a hit with those particular time and ADC values is a signal hit. For a cell with multiple hits, we examine the time and ADC values of each hit and select the hit that corresponds to the highest density (or likelihood) in the 2D distribution. By selecting the hit with the highest likelihood, we can achieve reasonable prediction accuracy.
For comparison, we also tested a simpler method that selects the first hit in each cell. In the initial evaluation, both methods demonstrated roughly comparable performance, likely due to the low hit multiplicity in each cell. A more detailed study of comparison will follow in this paper.
@@ 3.5.2 Converting Hits to Pixels
After reducing the dimensionality by selecting only one hit per cell, each pixel corresponds to a single hit. Thus, only three color channels are needed for pixels. The first channel represents the hit timing, the second channel represents the ADC value, and the third channel represents $\Delta\phi$, the azimuthal angle between the CTH trigger counter and each cell.
The three types of input values can be converted to color values using simple linear scaling. However, one feature in the distribution of signal hits suggests that scaling with a limited input value range might be beneficial for maintaining precision with a limited number of bits for color values. In the simulation, as shown in Figure 8, most signal electron hits fall within the range (--50.0 ns, 400 ns) for time information and (0, 700) for the ADC information. In the limited scaling scheme, hits within these ranges are linearly scaled, while those outside are assigned capped values.
When mapping hit cells to pixels, some pixels correspond to non-existing cells, while some correspond to cells with no hits. To distinguish these special pixels from others, black is reserved for pixels without corresponding cells, while white is used for pixels representing empty cells. Narrow gaps in the color value range are maintained around both black and white to ensure these special cases are clearly distinguishable from regular hit data.
Figure 8: Signal and background hit distributions for hit timing (left) and hit ADC (right) are shown with logarithmic-scaled y-axis. Signal hits cluster within specific ranges, while background hits are more uniformly distributed.
@@ 3.5.3 Arranging Pixels on Canvas
In the ”normal” format, the detector’s 2-dimensional plane—more specifically, a projection of cells at the CDC readout plane—is treated as an image, using (x, y) coordinates to define the plane on the CDC section. The cell size is approximately 16.8 mm wide and 16.0 mm high, with the radius of the CDC readout at 818 mm [3]. To represent each detector cell as a single pixel, the plane is discretized into 128 bins along both the x and y axes, resulting in a 128$\times$128 pixel image. This results in a grid size of 12.8 mm per bin (1636 mm / 128), which is smaller than the actual cell size. Although the cells are arranged along the section of a cylinder, requiring a finer grid when using Cartesian coordinates, this discretization is essential for ensuring each cell is accurately represented by a single pixel. However, it is important to note that approximately 70% of this image area remains unoccupied(black area of Figure 9(a)), rendering this format inefficient for neural network training.
@@
Figure 9: (a),(b) The input image and label of the normal format. (c),(d) The input image and label of the reduced format. Each input image has 3 channels: red for hit timing, green for ADC of hits, and blue for $\Delta\phi$ of hits. White pixels of input image indicate cells with no hits, while black pixels represent areas not corresponding to any cells in the CDC detector. The label images are gray scale, with white pixels denoting cells with true signal hits.
@@ 4.1.1 U-Net
EfficientNet, developed through neural architecture searches [20, 21], optimizes both accuracy and efficiency by scaling up depth, width, and resolution in a balanced manner. This model builds on the success of earlier architectures like ResNet [4], incorporating advancements that allow it to achieve superior performance with fewer parameters and reduced computational cost. EfficientNet comes in a family of models ranging from EfficientNet-B0 to EfficientNet-B7, each with an increasing number of parameters and layers. This scalability allows us to evaluate different model capacities by simply changing the model index, facilitating a more systematic investigation of the model’s performance on our task.
@@ 5 Implementation Details
Our preprocessing methods include options such as linear and limited scaling for the scaling method, and the first hit and 2D distribution for handling multi-hit events. By default, we use linear scaling for the scaling method and the 2D distribution method for multi-hit handling.
@@ 6 Results and Analysis
This metric provide a comprehensive evaluation of the model’s ability to accurately identify positive and negative cases, a critical aspect for addressing the requirements of COMET tracking task. And it is said that to obtain the necessary momentum resolution in COMET, this Purity-Retention curve must pass outside (0.9,0.9), so we will discuss whether such a learning result can be obtained.
@@ 6.1 Model Performance on Signal Detection
Our model achieved the target score, surpassing both purity and retention metrics of 0.9 simultaneously, as illustrated in Figure 11. We can visually confirm our model successfully picks up target signal information from noisy input as shown in Figure 12. From Figure 11, We identified two key factors essential for achieving the target score:
@@ 6.2 Effect of Scaling Model Parameters
Another reason for the reduced sensitivity of parameter scaling at larger image sizes is the nature of the current background hits, which are generated with the resampling method (Sec.3.2) and can be considered relatively easy to classify compared to real cases. Consequently, the models can achieve high accuracy even with fewer parameters when the image size is sufficient to capture the necessary details.
@@ 6.3 Handling Multiple Hits
The Purity-Retention curves in Figure 15 compare the performance of two preprocessing methods for handling multiple hits within the same cell: the naive 1st-hit method and the 2D distribution method. The left curve shows the overall Purity-Retention relationship, while the right zooms in on the high-purity region (close to 1.0). The 2D distribution method (green dashed line) consistently outperforms the 1st-hit method (blue solid line) across almost the entire range of retention values, particularly in the high-retention region, where it maintains higher purity levels compared to the 1st-hit method.
The superior performance of the 2D distribution method can be attributed to its enhanced signal hit coverage, as highlighted in Table 3. The table shows that the 1st-hit method has lower overall signal coverage (87%) and significantly lower signal coverage for multi-hit cells (74%). In contrast, the 2D distribution method achieves a higher overall signal coverage (97%) and better coverage for multi-hit cells (94%). This improved coverage translates directly into higher purity at comparable retention levels, as demonstrated in the Purity-Retention curves. Although the techniques we explored are relatively basic, their promising results warrant further exploration in future studies.
Figure 15: Purity-Retention curves for two methods. The left plot shows the full-scale view of the Purity-Retention curves for the 1st_hit (solid blue line) and the highest_hit (dashed green line). The 1st_hit refers to using the first hit method for multiple hits handling, while the 2D_histogram refers to selecting the highest value with the 2D distribution method to improve model performance. The right plot is a zoomed-in view focusing on the high-purity and high-retention region, indicated by the yellow area. This highlighted area represents where both purity and retention are above 0.9, which is our target.
Table 3: Comparison of Signal Coverage between the 1st Hit and 2D distribution methods. This statistics is calculated with 8000 events from the training data. The 2D distribution method shows better coverage than the 1st hit method, which explains the superior performance of the 2D distribution in the Purity-Retention curve.